OpenStudy (anonymous):

Solve: 3(x + 1) + 6 = 33

OpenStudy (anonymous):

3(x+1)+6=33 3(x+1)=27 x+1=9 x=8

OpenStudy (anonymous):

first you should Distribute 3 to x and 1.

OpenStudy (anonymous):

3x+3+6=33 3x+9=33 3x=33-9 3x=24 x=24/3 x=8 answer